# Turn Every Application into an Agent: Towards Efficient Human-Agent-Computer Interaction with API-First LLM-Based Agents

### v1

@@ 3.3.3. Skill Validation
Static Validation: This validation method utilizes structural method to verify the skill code, including checking whether the skill’s parameters contain the mandatory parameters (such as the executor instance and args list), whether the methods and properties of the executor are correctly invoked in the code, and whether any non-existent skills are imported when reusing the skill.
Dynamic Validation: This validation method evaluates a skill’s performance in the app environment with the help of two agents: ValidatorAgent and EvaluateAgent. When a skill is submitted for validation, ValidatorAgent firstly proposes a task based on the skill’s functional description. It then executes the skill within the same app environment used during the initial skill exploration and records the responses and environmental changes during execution. Upon the completion, the EvaluateAgent determines whether the skill has successfully completed the task or not.
4. Feasibility Study
@@
To validate the usability and effectiveness of AXIS framework, we conducted a feasibility study. We first use AXIS to explore Microsoft Office Word and discover 73 skills. Then we extracted 50 tasks from the wikihow 333https://www.wikihow.com/Use-Microsoft-Word page ”Use Microsoft Word” and the official Microsoft Word website 444https://support.microsoft.com/en-us/word. These tasks were executed using both AXIS and UI Agent, and the results were analyzed and compared.
4.1. SKill Exploration
@@
Before the exploration, AXIS was provided with 6 basic actions, as shown in Table 2. Then, 347 seed files were used for the AXIS to explore. After the exploration, AXIS discovered 73 skills with different hierarchies. Majority of the skills (44) discovered have a skill hierarchy 1. The rest is composed of 24 skills with hierarchy 2, 3 skills with hierarchy 3, and 2 skills with hierarchy 4. Table 3 displays several successfully validated skills discovered during the exploration process.
| Name | Description | Example |
| set_edit_text | The function to Set the edit text of the control element, can use to input content on Edit type controls. | set_edit_text(executor, args_dict=”control_id”:’119’, ”control_name”:”Edit”, ’text’:”hi there”) |
| select_text | A function to select the text with the specified text content. | select_text(executor, args_dict=”text”:”hello”) |
| select_table | A function to select the table with the specified number. | select_table(executor, args_dict=”number”:1) |
| type_keys | A function to Type in keys on control item.Used to enter shortcuts and so on. | type_keys(executor, args_dict=”control_id”:’119’, ”control_name”:”Edit”, ”text”: ”VK_CONTROL down”, ”newline”: False) |
| click_input | A function to Click the control element.Usually be used to switch to different ribbon,click the buttons in menu. | click_input(executor, args_dict=”control_id”:”12”, ”control_name”:”Border”, ’button’:”left”,’double’:False) |
| wheel_mouse_input | A function for Wheel mouse input on the control element. | wheel_mouse_input(executor, args_dict=”control_id”:”12”, ’wheel_dist’:-20) |
| Hierarchy | Name | Description | Example |
| 2 | align_text | The function aligns the text in a Microsoft Word document. It first selects the text, then applies the desired alignment (left, center, right, justify) using the Word API. | align_text(executor, args_dict=”text”: ”hello”, ”alignment”: ”center”) |
| 3 | apply_text_style | A function to edit a text with specified text, font size, font name. The title is set in the center. | apply_text_style(executor, args_dict=”text”:”Hello”, ”font_name”:”Arial”, ”font_size”:13) |
4.2. Task Completion
The 50 Word-related tasks collected above were used to test and compare the performance of UI Agent and AXIS with the explored skills. In our experiment, we choose UFO (Zhang et al., 2024a) as the representative of UI Agent considering its good performance on word tasks. The results are presented in Table 4, which includes the average time taken to complete different tasks, the success rates, the average number of steps per task, and the corresponding costs of LLM backend (GPT-4o, version 20240513) for both agents.
@@ 4.2. Task Completion
In terms of execution time, AXIS significantly outperforms the UI Agent, with an average task completion time of 29.9 seconds compared to 59.5 seconds for the UI Agent. This result shows that AXIS is nearly twice as fast as the UI Agent. AXIS also achieves a higher success rate in completing the tasks. Moreover, thanks to the abstraction and integration of basic actions into higher-level skills, AXIS could complete tasks with fewer steps on average, which also incurs lower costs compared to the UI Agent UFO.
| Metric | UI Agent | AXIS | Pairwise Significance |
| Time(s) | 59.5 | 29.9 | u¿a (p ¡ 0.001) |
| Success Rate(%) | 52.0 | 84.0 | u¡a (p ¡ 0.001) |
| Steps | 3.2 | 2.0 | u¿a (p ¡ 0.01) |
| Cost($) | 0.4 | 0.2 | u¿a (p ¡ 0.001) |
To better understand the reasons behind AXIS’s higher efficiency, we analyzed the number of UI and API-type actions invoked by AXIS and the UI Agent during task execution and recorded the proportion of API and Advanced API (defined as skills with a hierarchy level of 2 or higher) calls made by both methods. As shown in Table 5, AXIS invoked significantly fewer UI actions compared to the UI Agent during task execution. Notably, the total number of UI actions performed by AXIS across all tasks was greater than the number of invoked API actions. Upon inspection, this was found that AXIS tends to use a single and integrated API skill to complete a whole task, resulting a low overall API actions count. We further calculated the API usage rate and the proportion of Advanced API usage among the API actions. The data shows that AXIS’s proportion of API actions reached 55.7% with a 23.1% usage rate of advanced API. In contrast, the API usage rate of UI Agent is only 8.1%. Based on the above results, we conclude that AXIS indeed adopts an API-first approach and tends to use skills to complete tasks when the matching skills are available. And the integration of skills into actions also significantly contributes to the increased efficiency of AXIS.
@@
5. User Study
We carry out extensive user experiment to evaluate the performance of AXIS. The experiment and the associated evaluation metrics were designed to explore the following research questions (RQs) regarding the roles of LLM-based agents in work and life scenarios:
RQ1: Does LLM-based agent lower the cognitive load of the users and make them have less effort to learn?
RQ2: Does LLM-based agent enhance the efficiency of users?
@@ 5. User Study
In our user experiment, participants were asked to complete specified tasks within an application through three methods: manually, with the assistance of a UI Agent, and with the assistance of AXIS and recorded the entire process. Microsoft Word is chosen as the experimental application considering its popularity in our daily work and life as well as the rich API documentations (Microsoft365, 2024b)). Motivated by the RQs, we set three objectives for the user experiment:
(1) To evaluate the cognitive load on participants when completing tasks using different methods. (2) To compare the efficiency and reliability of task completion across the three methods. (3) To assess user preferences regarding the use of different Agents. This study is approved by the Institutional Review Board (IRB) of Peking University.
5.1. Experiment Procedure
@@
We provided users with a simple web interface during the formal experiment, which consisted of two stages. In Stage 1, participants received a pre-printed task list including both L1 and L2 tasks. Based on the task ID displayed on the webpage, participants were asked to read the task requirements, click the ”start” button, complete the task in the automatically opened Word document, and click ”Finish.” upon task completion. In Stage 2, participants were instructed to use both the UI Agent and AXIS to assist them in completing the Word tasks. The corresponding webpage were featured with both input fields and buttons for activating the two Agents. The participants need to enter task description to command the Agents to complete the tasks. Throughout the formal experiment, all task execution processes were recorded for subsequent analysis. After completing all assigned tasks manually or with the assistant of agents, four different post-task questionnaires were displayed on the experimental webpage to survey users’ subjective experiences.
5.2. Participants

### v2

@@ 4.3 Stage III: Skill Validation
Static Validation. Static validation utilizes structural checks to verify the compatibility of the skill code with the skill executor. It examines whether the skill’s parameters contain the mandatory parameters (such as the executor instance and args list), whether the methods and properties of the executor are correctly invoked in the code, and whether any non-existent skills are imported when reusing the skill. Skills that do not meet these formal requirements are returned for revision.
Dynamic Validation. Dynamic validation evaluates the performance of a skill in a real environment. It consists of a validator and an evaluator. The validator generates various input parameters to test the skill’s generalizability, and the evaluator checks whether the test is successfully completed by examining the final state. The specific prompts are detailed in Appendix LABEL:sec:skillval-prompts, with detailed description of the validator and evaluator in Appendix LABEL:appendix:dynamicvalidator.
5 Feasibility Study
@@
To validate the usability and effectiveness of the AXIS framework, we conduct a feasibility study. We first use AXIS to explore Microsoft Word and discover 73 skills. Then we extract 50 tasks from the wikihow 444https://www.wikihow.com/Use-Microsoft-Word page “Use Microsoft Word” and the official Microsoft Word website 555https://support.microsoft.com/en-us/word. These tasks were executed using both AXIS and UI Agent, and the results were analyzed and compared. AXIS also enable turn an application into an agent by simply wrapping the application with an API set and adopting a simpler UI design suitable for HACI, which is presented in Appendix LABEL:sec:extensive.
5.1 SKill Exploration
@@
Before the exploration, AXIS is provided with the initial skill library which composed of 6 basic actions as shown in Table 7. Then, 347 seed files are used for AXIS to explore. After the exploration, AXIS discovers 73 skills with different hierarchies. Majority of the skills (44) discovered have a skill hierarchy 1. The rest is composed of 24 skills with hierarchy 2, 3 skills with hierarchy 3, and 2 skills with hierarchy 4. Table LABEL:tab:explored displays several successfully validated skills discovered during the exploration process.
5.2 Task Completion
We evaluate UI Agent (represented by UFO Zhang et al. (2024a) for its superior Word task performance) and AXIS on 50 Word-related tasks using explored skills. Table 1 presents comparative results, including completion time, success rates, step counts, and LLM backend costs (GPT-4o, version 20240513) for both agents.
@@ 5.2 Task Completion
In terms of execution time, AXIS demonstrates superior performance, completing tasks in 29.9 seconds on average - twice as fast as UI Agent’s 59.5 seconds. It also achieves higher success rates and, through its skill-based abstraction, requires fewer steps per task, resulting in lower costs compared to the UI Agent UFO.
| Metric | UI Agent | AXIS | P - Significance |
| Time(s) | 59.5 | 29.9 | u>a (p < 0.001) |
| Success Rate(%) | 52.0 | 84.0 | u<a (p < 0.001) |
| Steps | 3.2 | 2.0 | u>a (p < 0.01) |
| Cost($) | 0.4 | 0.2 | u>a (p < 0.001) |
To investigate AXIS’s efficiency, we analyze action types and API usage patterns. As shown in Table 2, AXIS employs significantly fewer UI actions compared to the UI Agent, while utilizing more API and Advanced API calls (skills with hierarchy level 2). Analysis reveals AXIS predominantly employs integrated API skills for task completion, yielding more API actions (55.7% usage rate, including 23.1% advanced APIs) compared to UI Agent’s 8.1%. This demonstrates AXIS’s API-first approach, leveraging available skills for efficient task execution through skill-action integration.
@@
6 User Study
We conduct an extensive user experiment to evaluate the performance of AXIS. The experiment and evaluation metrics are designed to explore the following research questions (RQs) on the role of LLM-based agents in work and daily life scenarios:
RQ1: Does the LLM-based agent lower the cognitive load of the users and make them have less effort to learn?
RQ2: Does the LLM-based agent enhance the efficiency of users?
@@ 6 User Study
In our user experiment, participants are asked to complete specified tasks within an application through three methods: manually, with the assistance of a UI Agent, and with the assistance of AXIS. The entire process is recorded. Microsoft Word is chosen as the experimental application considering its popularity in our daily work and life as well as the rich API documentations Microsoft365 (2024b)). Thus five tasks of Word are sampled from both official Word documentation and GPT-generated results, divided into two difficulty levels: low difficulty (L1) and high difficulty (L2). Motivated by the RQs, we set three objectives for the user experiment: (1) To evaluate the cognitive load on participants when completing tasks using different methods. (2) To compare the efficiency and reliability of task completion across the three methods. (3) To assess user preferences regarding the use of different Agents. This study is approved by the Institutional Review Board (IRB) of University.
The details of experiment procedure and participants recruitment are shown in Appendix LABEL:sec:user-study.
| Metric | Task Level | Manual | UI Agent | AXIS | Pairwise Significance |
| Time(s) | L1 L2 | 61.8 167.6 | 104.6 155.5 | 18.2 57.1 | L1: m<u (p < 0.001) L1, L2: a<m (p < 0.001) L1, L2: a<u (p < 0.001) |
| Success Rate(%) | L1 L2 | 100.0 97.5 | 75.0 45.0 | 98.3 95.0 | L1, L2: m>u (p < 0.001) L1, L2: a>u (p < 0.001) |
| Metric | Task Level | UI Agent | AXIS | Pairwise Significance |
| steps | L1 L2 | 6.4 11.1 | 1.0 4.2 | L1: a<u (p < 0.001) L2: a<u (p < 0.001) |
| cost($) | L1 L2 | 0.6 0.9 | 0.07 0.3 | L1: a<u (p < 0.001) L2: a<u (p < 0.001) |
6.1 Experimental Metrics
We collect both subjective and objective metrics in our experiments to evaluate the performance and user experience of different methods.
Subjective Metrics. As detailed in Appendix LABEL:sec:user-study, we conduct four post-task questionnaires (manual and Agent-assisted L1/L2 tasks) using NASA-TLX Hart (1988) metrics: Mental, Physical, and Temporal Demand; Performance; Frustration; and Effort, supplemented by a learning effort metric. Lower scores indicate reduced cognitive load, improved success perception, and decreased frustration/effort. For Agent-assisted tasks (Questionnaires 3-4), additional metrics include Agent fluency/reliability, UI dependency, decision consistency (Agent-user alignment), and perceived completion speed.
Objective Metrics. For objective metrics, we maintain comprehensive experimental logs, including screen recordings of manual and Agent-assisted tasks, decision-making processes, UI interaction paths, task completion time and success rates, LLM backend costs (GPT-4, version 20240513) for Agent operations, and UI dependency measurements for both the UI Agent and AXIS.
| Metric | Task Level | Manual | Agents | Pairwise Significance |
| Mental Demand (0-100) | L1 L2 | 21.3 70.0 | 2.5 7.5 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Physical Demand (0-100) | L1 L2 | 31.3 57.5 | 5.0 6.3 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Temporal Demand (0-100) | L1 L2 | 52.5 37.5 | 28.8 35.0 | L1: m>a (p < 0.05) L2: - |
| Performance (0-100) | L1 L2 | 21.2 47.5 | 21.2 26.2 | L1: - L2: m>a (p < 0.05) |
| Frustration Level (0-100) | L1 L2 | 31.3 62.5 | 7.5 10.0 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Completion Effort (0-100) | L1 L2 | 12.5 35.0 | 17.5 13.8 | L1: - L2: m>a (p < 0.01) |
6.2 Results
Our analysis explores three key aspects: (1) cognitive load reduction, (2) task efficiency of agents, and (3) user preferences between UI agents and AXIS.
Cognitive Load. To assess cognitive load reduction by LLM-based Agents, we analyze NASA-TLX and learning effort scores (Table 5, Figure 3). Results show higher scores for L2 tasks across dimensions, validating our task difficulty classification. Agent-based methods significantly outperform manual approaches in reducing mental/physical demand and frustration, particularly for complex L2 tasks (p<0.05). While L1 task performance differences were minimal, agents notably enhanced users’ success perception in L2 tasks. Analysis reveals consistent user experiences across task complexities (Figure 3 (b)) and significant learning effort reduction, especially for difficult tasks (Figure 3 (c)). These findings answer RQ1: LLM-based agents effectively reduce cognitive load and learning effort, particularly for complex tasks.
Efficiency and Reliability. Our evaluation compares manual, UI Agent, and AXIS approaches using completion time, success rate, steps, and costs (summarized in Tables 3 and 4). AXIS demonstrates superior time efficiency, significantly outperforming both methods (p < 0.001), particularly in complex L2 tasks. While manual methods achieve highest accuracy, AXIS maintain near-human performance, contrasting with UI Agent’s lower accuracy due to UI element positioning and visibility issues. Notably, UI Agents require substantially more steps for deeply nested L2 tasks, whereas AXIS’s API-driven approach enable streamlined execution with fewer steps and lower costs. This answers RQ2: while UI Agents offer modest efficiency gains for complex tasks with reliability challenges, AXIS consistently improves human efficiency with superior reliability.
Affinity Preference. To compare user experiences between UI agents and AXIS, we conduct a subjective evaluation across five aspects (Figure 4). Participants consistently preferr AXIS for its perceived speed, fluency, and reliability across both L1 and L2 tasks. While AXIS’s highly encapsulated API led to less human-like decision-making in simple tasks (L1), its decisions aligned better with human reasoning in complex tasks (L2). AXIS notably improves user’s experienct by reducing UI dependency compared to UI agents’ frequent interface interactions. These findings answer RQ3: AXIS provides superior efficiency, smoothness, and reliability compared to UI agents, with increasing user preference as task complexity grows. User feedback highlights AXIS’s API-first approach as advantageous over UI agents’ intrusive behaviors, offering better control and experience.
7 Conclusion
We present AXIS, a framework that enhances human-agent-computer interaction (HACI) by prioritizing API calls over UI interactions to reduce inefficiencies and cognitive burdens in complex tasks with multimodal large language models (MLLMs). Experiments with Microsoft Word show that AXIS cuts task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining human-level accuracy. These results demonstrate the potential of API-first LLM-based agents to streamline interactions, reduce latency, and improve reliability, offering a novel approach to faster, more efficient task execution.
8 Ethical Statement
All used datasets collected internally or obtained from external sources, ensuring no infringement on individual or organizational rights. User study participants volunteered and were compensated.
9 Limitations
While AXIS effectively mines APIs and enables efficient human-agent-computer interaction (HACI), further optimization is needed to achieve an Agent OS. Currently, AXIS primarily relies on Python-based APIs, limiting support for applications without native Python interfaces. Additionally, its exploration process requires improvements in stability and efficiency. Future work should focus on developing unified action interfaces to extend HACI to more applications and operating systems while enhancing the framework’s performance and efficiency.
Appendix A Interfaces in Environment
Below are the details of two interfaces: state() and step() to return the environment state and execute agent actions, respectively.
state(): Returns the state of the environment, including detailed information on the current elements within the environment. The environment state encompasses key UI information such as control positions, control types, and selection status, which is consistent with the definition in Zhang et al. (2024a). For applications that can be unpacked, the unpacked XML content is also included as part of the state.
step(): Incorporates a skill executor that allows agents to perform operations within the environment by executing skills. Upon completion, this interface returns the results of these operations.
Appendix B Skills
B.1 Skill Component
Each skill consists of three main components:
Skill Code: A piece of code structured to be compatible with the skill executor. Skill code includes a uniform set of parameters and adheres to the standard PEP 257 documentation. The initial set of skills is generated by restructuring the fundamental APIs from the application provider, and based on these initial skills, AXIS can explore and develop additional new skills.
Description: A description of the skill’s functionality, which assists the LLM in selecting and invoking appropriate skills during task execution.
Usage Example: One or more code examples including typical parameters associated with the code and description. These examples help the LLM in correctly formatting parameters when invoking the skill.
Skill Executor. As discussed above, our application environment incorporates a step() interface to facilitate the interaction between agents and the environment. This interface also hosts the skill executor responsible for executing the skill generated or selected by the agents. The skill executor keeps caching of application documents and simultaneously supports multiple functionalities including locating application controls, invoking methods on those controls, and calling application APIs (independent of controls), to enable the UI actions and API actions in the same time and serve as an efficient foundation for skill-driven operations.
B.2 Skill Type
The skills in AXIS can be categorized into five types based on the types of the code fragments: Atomic UI Skill, Atomic API Skill, Composite UI Skill, Composite API Skill, and API-UI Hybrid Skill. There are some examples in Table 6.
| Type | Description | Example | Feature coverage |
| Composite API skill | Composed of multiple atomic API actions or composite API actions. This type of skills often represents a higher-level combination of functions. | insert_header_footer | insert header and footer with specified contents by API, which is equal to sequential UI actions "Insert->Header->footer edit->Footer->footer edit". |
B.3 Initial Skill Repository
As the foundation for the interaction between the agent and the environment, before exploring and mining skills, we pre-defined some basic actions as the initial skill library shown in Table 7. These actions are derived from UFO Zhang et al. (2024a).
| Name | Description | Example |
| set_edit_text | The function to Set the edit text of the control element, can use to input content on Edit type controls. | set_edit_text(executor, args_dict="control_id":’119’, "control_name":"Edit", ’text’:"hi there") |
| select_text | A function to select the text with the specified text content. | select_text(executor, args_dict="text":"hello") |
| select_table | A function to select the table with the specified number. | select_table(executor, args_dict="number":1) |
| type_keys | A function to Type in keys on control item.Used to enter shortcuts and so on. | type_keys(executor, args_dict="control_id":’119’, "control_name":"Edit", "text": "VK_CONTROL down", "newline": False) |
| click_input | A function to Click the control element.Usually be used to switch to different ribbon,click the buttons in menu. | click_input(executor, args_dict="control_id":"12", "control_name":"Border", ’button’:"left",’double’:False) |
| wheel_mouse_input | A function for Wheel mouse input on the control element. | wheel_mouse_input(executor, args_dict="control_id":"12", ’wheel_dist’:-20) |
B.4 Skill Translation Example
The following is a specific example regarding the generation and translation of skills in Table LABEL:tab:skill-trans.
| Generated Skill by Generator Agent | Translated Skill by Translator Agent |
| ⬇ code: |- from select_text import select_text def hightlight_text(executor, args_dict: dict): """ Highlight the text. :param executor: The executor object. :param args_dict: The arguments of the highlight text method in a dict - text: the text to highlight - home_control_id: the id of the Home ribbon. :return: The result of the action.True or False. """ try: select_text(executor, args_dict) home_control_id = args_dict.get("home_control_id", -1) home_ribbon_control = executor.get_target_control_by_args( control_args={"control_id":home_control_id}) executor.atomic_execution( control=home_ribbon_control, method_name="click_input", args={ "button": "left", "double": False } ) hightlight_button = executor.get_target_control_by_uuid( "uuid_control.json", "96d5a516-35af-459b-b1c4-34aced6acb0b") executor.atomic_execution( control=hightlight_button, method_name="click_input", args={ "button": "left", "double": False } ) except Exception as e: print(f"error:{str(e)}") return False return True description: |- A function to highlight the text. Note: the args_dict should include "home_control_id" field of "Home" ribbon control id and "text" field of the target text to highlight. example: |- highlight_text(executor, args_dict={"home_control_id": 27, "text": "Hello"}) | ⬇ code: |- from select_text import select_text def highlight_text(executor, args_dict: dict): """ Highlight the text. :param executor: The executor object. :param args_dict: The arguments of the highlight text method in a dict - text: the text to highlight :return: The result of the action. True or False. Example: >>> highlight_text(executor, args_dict={"text": "hello"}) """ try: select_text(executor, args_dict) executor.app.Selection.Range.HighlightColorIndex = 7 # 7 represents yellow in Word except Exception as e: print(f"error: {str(e)}") return False return True description: |- The function is to highlight the text in the word document. It first finds the text to be highlighted, then selects the text, and finally highlights the text with yellow color. example: |- highlight_text(executor, args_dict={"text": "hello"}) |
B.5 Explored Skill Examples
Here are some explored skill examples of different hierarchies shown in Table LABEL:tab:explored.
| Hierarchy | Name | Description | Example |
| 2 | align_text | The function aligns the text in a Microsoft Word document. It first selects the text, then applies the desired alignment (left, center, right, justify) using the Word API. | align_text(executor, args_dict="text": "hello", "alignment": "center") |
| 3 | apply_text_style | A function to edit a text with specified text, font size, font name. The title is set in the center. | apply_text_style(executor, args_dict="text":"Hello", "font_name":"Arial", "font_size":13) |
B.6 Skill Dynamic Validator
Dynamic validation evaluates the performance of a skill in actual tasks and environments through two specialized agents as detailed below:
Validator Agent: When a skill is submitted for validation, this agent analyzes the skill’s code, usage examples, and description to generate appropriate test tasks. Create test parameters consistent with the skill’s usage examples and execute the skill in randomly generated Word documents to assess its functionality.
Evaluator Agent: After the Validator Agent executes the target skill, the Evaluator Agent examines the execution logs, document modifications, and final state of the Word document. Evaluate whether the task was successfully completed according to the intended functionality. Only skills that pass this evaluation are eventually added to the skill library.
Appendix C User Study
C.1 Experiment Procedure
@@
We provided users with a simple web interface during the formal experiment, which consisted of two stages. In Stage 1, participants received a pre-printed task list including both L1 and L2 tasks. Based on the task ID displayed on the webpage, participants were asked to read the task requirements, click the "start" button, complete the task in the automatically opened Word document, and click "Finish." upon task completion. In Stage 2, participants were instructed to use both the UI Agent and AXIS to assist them in completing the Word tasks. The corresponding webpage were featured with both input fields and buttons for activating the two Agents. The participants need to enter task description to command the Agents to complete the tasks. Throughout the formal experiment, all task execution processes were recorded for subsequent analysis. After completing all assigned tasks manually or with the assistant of agents, four different post-task questionnaires were displayed on the experimental webpage to survey users’ subjective experiences.
C.2 Participants Recruitment